# CERN-TH-2025-012, IPPP/25/09, ZU-TH 11/25 NNLOJET: a parton-level event generator for jet cross sections at NNLO QCD accuracy

### v1

@@ 1 Introduction
Many precision observables are derived from hadronic final states, typically in the form of identified jets, which can be related to the production of quarks and gluons (partons) in hard scattering processes [1, 2]. These jet observables provide clean probes of the dynamics of the strong interaction (quantum chromodynamics, QCD) at electron–positron colliders and they affect all types of precision measurements at lepton–hadron or hadron–hadron colliders due to the partonic structure of the colliding hadrons. The theoretical description of these observables must take account of the detailed definition of the final state that is applied in the experimental measurement by incorporating the jet clustering algorithm and any fiducial cuts and isolation requirements on the final-state objects. The resulting predictions for fiducial cross sections can then be compared directly to the experimental data, thereby they stress-test the Standard Model and refine the determination of the parameters (Standard Model couplings and masses, parton distribution functions) that enter the theoretical predictions.
@@
Theoretical predictions rely on a perturbation theory expansion in the coupling constants. In the case of hadron-collider processes and jet observables, QCD corrections are particularly important due to the large value of the strong coupling constant compared to the other Standard Model interactions. At higher orders in perturbation theory, a Born-level process with given final-state multiplicity receives two types of contributions: virtual corrections from closed particle loops in the underlying scattering amplitudes and real-radiation corrections from final states with higher multiplicity. At next-to-leading order (NLO), these correspond to one-loop virtual corrections (V) and single real emission corrections (R). With increasing orders, corrections can be of mixed type, with next-to-next-to-leading order (NNLO) consisting of two-loop double virtual (VV), mixed one-loop single real emission (RV) and tree-level double-real emission corrections (RR) and so on. Each individual subprocess contribution contains infrared singularities arising from soft or collinear momentum configurations in loop integrals and real-radiation processes; only the sum of all subprocesses contributing to a well-defined (infrared-safe, [3]) final state is finite and physically meaningful. These infrared singularities are explicit (as poles in the dimensional regularization parameter) only in the virtual loop amplitudes [4, 5, 6, 7], while leading to divergent phase-space integrals in the real-radiation processes. To obtain numerically finite predictions from real-radiation corrections, their infrared singularity structure must be made explicit.
@@
For this task, various infrared subtraction methods have been proposed and implemented for NLO calculations [8, 9, 10]. When combined with generic procedures for the evaluation of one-loop virtual amplitudes [11, 12, 13, 14] for any process, these methods have enabled the full automation of NLO corrections in QCD and the electroweak theory, becoming part of multi-purpose event simulation programs [15, 16, 17, 18, 19], which are complemented by the dedicated libraries for NLO processes in MCFM [20] and VBFNLO [21]. These perturbative fixed-order calculations are implemented as parton-level Monte Carlo event generation programs. Such a program contains the infrared-finite remainders of all parton-level subprocesses. These are evaluated over their respective phase spaces of different multiplicities, producing weighted parton-level events with full kinematic information that can be analysed according to the experimental fiducial cross section definition.
@@
Building upon the universal factorization of QCD real-radiation amplitudes in their infrared limits [22, 23, 24, 25, 26], several NNLO subtraction methods have been formulated [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. These methods have enabled the calculation of NNLO QCD corrections to a large number of collider processes [39] with up to three identified final-state objects. Such calculations are performed on a case-by-case basis in the framework of parton-level Monte Carlo event generation, typically requiring high-performance computing resources to yield stable and numerically reliable results. The resulting codes often require dedicated tuning and post-processing for each observable. They are usually not broadly available as open-source. Notable exceptions are EERAD3 [40, 41] for jet production in electron-positron annihilation, the MATRIX [42] and MiNNLOPS [43, 44] codes for processes with colour-neutral or heavy-quark final states, the NNLOCAL code [45] for colour-neutral final states and the NNLO-upgrade of the MCFM code [20, 46] for selected di-boson processes.
@@
In this work, we present the open-source release of the NNLOJET parton-level event generator code, which employs the antenna subtraction method [30, 47, 48] to compute the NNLO QCD corrections to several classes of jet production processes in hadron–hadron, lepton–hadron and electron–positron collisions. In the following, we review the antenna subtraction method in Sect. 2, describe the NNLOJET installation and general usage in Sect. 3 and the currently available processes in Sect. 4. The detailed usage of NNLOJET is documented through its runcard syntax and process workflow in Sections 5–6. The NNLOJET code is available on HEPForge: http://nnlojet.hepforge.org and is distributed under the GNU General Public License v3.0.
@@ 2 Antenna subtraction
The infrared singularities at NLO and NNLO are of two forms: explicit -poles from loop integrals in the virtual corrections [4] and implicit soft and/or collinear singularities in the real corrections [22, 23, 24, 25, 26], which become explicit only upon integration over the respective phase space. Consequently, it is not possible to implement equations (4) and (2) directly in a numerical code. To accomplish the cancellation of infrared singularities among the different contributions at NLO and NNLO, one introduces a subtraction procedure that extracts the implicit infrared singularities from real-radiation processes, makes them explicit by phase-space integration and combines them with the virtual corrections. The resulting expressions read
@@
NNLOJET implements the antenna subtraction method [30, 47, 48], which derives the subtraction terms starting from colour-ordered decompositions of all real-radiation contributions. The colour-ordering allows identifying each unresolved real-radiation configuration with a pair of hard radiator partons, which emit the unresolved partons. This configuration is then described by an antenna function consisting of the hard radiators and unresolved radiation in between them. Antenna functions were derived at NLO and NNLO for quark–antiquark [49], quark–gluon [50] and gluon–gluon [51] pairs as hard radiators. All quarks are treated as massless. The unintegrated subtraction terms are obtained from products of antenna functions and reduced matrix elements of lower partonic multiplicity, which are summed over all configurations of hard radiators and unresolved partons. The kinematics of the antenna functions and the reduced matrix elements are factorized from each other through a phase-space mapping [52, 47], such that the antenna function can be integrated over the antenna phase space of the hard radiator partons and the unresolved radiation while leaving the kinematics of the reduced matrix element unaffected. The integrated antenna functions enter the subtraction terms and where they establish the explicit cancellation of infrared -poles between real and virtual contributions. They were computed up to NNLO for all different configurations of the hard radiators: final–final [30], initial–final [53] and initial–initial [54, 55, 56].
With these ingredients, all terms in equation (2) can be implemented in the antenna subtraction formalism, as described in detail in [48]. For all processes available in NNLOJET, this implementation has been made on a process-by-process basis, usually starting from the implicit infrared singularity structure of the double-real-radiation process . A fully automated construction of the antenna subtraction terms in colour space is currently under development [57, 58].
@@ 3 The NNLOJET framework
The NNLOJET code is a parton-level event generator that provides the framework for the implementation of jet production processes to NNLO accuracy in QCD, using the antenna subtraction method. It contains the event generator infrastructure (Monte Carlo phase-space integration, event handling and analysis routines) and provides the unintegrated and integrated antenna functions and the phase-space mappings for all kinematic configurations. The NNLOJET phase-space integration is based on optimized parametrizations of the respective subprocess phase spaces (described in detail in [59]), which are then integrated using the adaptive Monte Carlo routine Vegas [60].
The implementation of processes in the NNLOJET framework requires the availability of the matrix elements for all RR, RV, and VV contributions, as well as the construction of the antenna subtraction terms. The subtraction terms have been validated by verifying their point-wise convergence to the respective subprocess matrix elements in all limits, as described in detail in [61, 57]. Some of the one-loop subprocess matrix elements were taken from the MCFM code [20], in particular for processes involving vector bosons [62]. The two-loop virtual corrections to all processes with four-point kinematics or beyond contain various special functions, in particular harmonic polylogarithms [63] and their generalizations. These are evaluated using HPLOG [64], TDHPL [65] and CHAPLIN [66], which are all included as part of the NNLOJET distribution. The evaluation of parton distributions and of the strong coupling constant is performed using the standard LHAPDF [67] interface.
@@ 3.1 External dependencies
The NNLOJET code is mainly written in modern Fortran with some modules, dependencies and driver files written in C++ and Python. Multi-threading capabilities are supported through OpenMP. The NNLOJET installation requires compilers for Fortran/C/C++, a Python 3 interpreter (minimum version 3.10), as well as CMake (minimum version 3.18) to be available on the system. In addition, LHAPDF 6 [67] must be present, ideally with the lhapdf-config executable searchable from the $PATH of the system. The user is responsible for downloading the necessary PDF sets to be used in the calculation.
@@ 4 Processes in NNLOJET
The processes included in the current release are described in the following. Details on their NNLOJET implementation and on validation tests are described in detail in the respective publications. In general, the following checks have been applied on all process: (1) point-tests [61, 57] and (2) technical-cut checks for RR and RV contributions; (3) finiteness of the RV and VV contributions.
@@
The pole tests are performed numerically for a large number of bulk phase space points on the RV process, where -poles of matrix elements, mass-factorization terms and subtraction terms are implemented explicitly. In the VV process, all integrated subtraction terms are combined with the mass-factorization terms and the pole terms extracted [4] from the double-virtual matrix element to obtain analytical pole cancellation and to define the finite remainder that is implemented numerically.
@@ 4.1 Jet production in electron–positron collisions
Jet final states in annihilation and related event shapes played a key role in establishing QCD as theory of the strong interaction and their theoretical study has been a major driver of developments in precision calculations in QCD. The cross section and related event shapes were the first jet observables to be computed to NNLO in QCD [68], originally implemented in the code EERAD3 [40] which is also based on antenna subtraction. The implementation of and in NNLOJET is described in [69]. It uses the matrix elements for vector-boson decay into partons [70, 71, 72, 73, 74, 75, 76], which account for the incoming lepton kinematics. The commonly used LEP-era event-shape variables and jet algorithms [77] are part of the generic NNLOJET infrastructure.
@@ 4.2 Jet production in electron–proton collisions
Jet production in deeply inelastic electron–proton scattering (DIS) allows to probe aspects of the proton structure that are difficult to assess in inclusive DIS and provides important precision tests of QCD. The implementation of up to NNLO QCD is described in [78, 79], including both neutral-current and charged-current DIS. It is based on crossings [75] of the matrix elements to DIS kinematics. Jets can be defined in the Breit frame or in the electron–proton collider frame and DIS event shapes [80] are part of the NNLOJET infrastructure.
@@ 4.3 Jet production in hadronic collisions
Jet production is the most ubiqutous QCD process at hadron colliders, probing the basic parton scattering reactions. A number of kinematic distributions can be derived from single-inclusive jet production and from dijet production . Their NNLOJET implementation is described in detail in [61, 81, 82, 83, 84, 85, 86, 87]. It is based on the relevant subprocess matrix elements up to NNLO [88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100]. The generic NNLOJET infrastructure allows selecting various different jet algorithms, performing scans over jet-resolution parameters and the setting of renormalization and factorization scales [101] either in a jet-based or event-based manner.
@@ 4.4 Vector-boson-plus-jet production
Precision studies of the electroweak interaction are commonly performed through massive vector-boson production, identified through their leptonic decay mode. Further kinematic information can be obtained by considering vector-boson-plus-jet final states. Observables derived from vector-boson and vector-boson-plus-jet production are widely used in precision measurements of electroweak parameters and of the proton structure. The NNLOJET implementation for , is described in [102, 103, 104, 105, 106, 107, 108] and for , in [109, 106]. The implementations are based on parton-level matrix elements up to NNLO, obtained as kinematic crossings [76] of the matrix elements. Decays into lepton pairs and the off-shell vector-boson propagator are included. It is also possible to obtain simplified predictions for on-shell vector-boson production in the narrow-width prescription.

### v2

@@ 1 Introduction
Many precision observables are derived from hadronic final states, typically in the form of identified jets, which can be related to the production of quarks and gluons (partons) in hard scattering processes [Ellis:1996mzs, Campbell:2017hsr]. These jet observables provide clean probes of the dynamics of the strong interaction (quantum chromodynamics, QCD) at electron–positron colliders and they affect all types of precision measurements at lepton–hadron or hadron–hadron colliders due to the partonic structure of the colliding hadrons. The theoretical description of these observables must take account of the detailed definition of the final state that is applied in the experimental measurement by incorporating the jet clustering algorithm and any fiducial cuts and isolation requirements on the final-state objects. The resulting predictions for fiducial cross sections can then be compared directly to the experimental data, thereby they stress-test the Standard Model and refine the determination of the parameters (Standard Model couplings and masses, parton distribution functions) that enter the theoretical predictions.
@@
Theoretical predictions rely on a perturbation theory expansion in the coupling constants. In the case of hadron-collider processes and jet observables, QCD corrections are particularly important due to the large value of the strong coupling constant s compared to the other Standard Model interactions. At higher orders in perturbation theory, a Born-level process with given final-state multiplicity receives two types of contributions: virtual corrections from closed particle loops in the underlying scattering amplitudes and real-radiation corrections from final states with higher multiplicity. At next-to-leading order (NLO), these correspond to one-loop virtual corrections (V) and single real emission corrections (R). With increasing orders, corrections can be of mixed type, with next-to-next-to-leading order (NNLO) consisting of two-loop double virtual (VV), mixed one-loop single real emission (RV) and tree-level double-real emission corrections (RR) and so on. Each individual subprocess contribution contains infrared singularities arising from soft or collinear momentum configurations in loop integrals and real-radiation processes; only the sum of all subprocesses contributing to a well-defined (infrared-safe, [Sterman:1977wj]) final state is finite and physically meaningful. These infrared singularities are explicit (as poles in the dimensional regularization parameter) only in the virtual loop amplitudes [Catani:1998bh, Becher:2009cu, Gardi:2009qi, Dixon:2009gx], while leading to divergent phase-space integrals in the real-radiation processes. To obtain numerically finite predictions from real-radiation corrections, their infrared singularity structure must be made explicit.
@@
For this task, various infrared subtraction methods have been proposed and implemented for NLO calculations [Giele:1993dj, Catani:1996vz, Frixione:1995ms]. When combined with generic procedures for the evaluation of one-loop virtual amplitudes [Ossola:2007ax, Ellis:2011cr, Cascioli:2011va, GoSam:2014iqq] for any process, these methods have enabled the full automation of NLO corrections in QCD and the electroweak theory, becoming part of multi-purpose event simulation programs [Alioli:2010xd, Alwall:2014hca, Bellm:2015jjp, Sherpa:2019gpd, Bierlich:2022pfr], which are complemented by the dedicated libraries for NLO processes in MCFM [Campbell:2019dru] and VBFNLO [Arnold:2008rz]. These perturbative fixed-order calculations are implemented as parton-level Monte Carlo event generation programs. Such a program contains the infrared-finite remainders of all parton-level subprocesses. These are evaluated over their respective phase spaces of different multiplicities, producing weighted parton-level events with full kinematic information that can be analysed according to the experimental fiducial cross section definition.
@@
Building upon the universal factorization of QCD real-radiation amplitudes in their infrared limits [Campbell:1997hg, Catani:1999ss, Catani:2000pi, Kosower:1999rx, Bern:1999ry], several NNLO subtraction methods have been formulated [Binoth:2004jv, Anastasiou:2003gr, Catani:2007vq, Gehrmann-DeRidder:2005btv, Czakon:2010td, Boughezal:2011jf, Boughezal:2015eha, Gaunt:2015pea, Cacciari:2015jma, Caola:2017dug, DelDuca:2016ily, Magnea:2020trj]. These methods have enabled the calculation of NNLO QCD corrections to a large number of collider processes [Heinrich:2020ybq] with up to three identified final-state objects. Such calculations are performed on a case-by-case basis in the framework of parton-level Monte Carlo event generation, typically requiring high-performance computing resources to yield stable and numerically reliable results. The resulting codes often require dedicated tuning and post-processing for each observable. They are usually not broadly available as open-source. Notable exceptions are EERAD3 [Gehrmann-DeRidder:2014hxk, Aveleira:2025svg] for jet production in electron-positron annihilation, the MATRIX [Grazzini:2017mhc] and MiNNLOPS [Monni:2019whf, Monni:2020nks] codes for processes with colour-neutral or heavy-quark final states, and the NNLO-upgrade of the MCFM code [Campbell:2019dru, Campbell:2022gdq] for selected di-boson processes, as well as a range of codes each covering individual processes at NNLO accuracy [Catani:2007vq, Anastasiou:2003yy, Catani:2009sm, Gavin:2010az, Catani:2011qz, Cacciari:2015jma, Camarda:2019zyx, DelDuca:2024ovc, Sherpa:2024mfk].
@@
In this work, we present the open-source release of the NNLOJET parton-level event generator code, which employs the antenna subtraction method [Gehrmann-DeRidder:2005btv, Daleo:2006xa, Currie:2013vh] to compute the NNLO QCD corrections to several classes of jet production processes in hadron–hadron, lepton–hadron and electron–positron collisions. In the following, we review the antenna subtraction method in Sect. 2, describe the NNLOJET installation and general usage in Sect. 3 and the currently available processes in Sect. 4. The detailed usage of NNLOJET is documented through its runcard syntax and process workflow in Sections 5–6. The NNLOJET code is available on HEPForge: http://nnlojet.hepforge.org and is distributed under the GNU General Public License v3.0.
@@ 2 Antenna subtraction
The infrared singularities at NLO and NNLO are of two forms: explicit -poles from loop integrals in the virtual corrections [Catani:1998bh] and implicit soft and/or collinear singularities in the real corrections [Campbell:1997hg, Catani:1999ss, Catani:2000pi, Kosower:1999rx, Bern:1999ry], which become explicit only upon integration over the respective phase space. Consequently, it is not possible to implement equations (4) and (2) directly in a numerical code. To accomplish the cancellation of infrared singularities among the different contributions at NLO and NNLO, one introduces a subtraction procedure that extracts the implicit infrared singularities from real-radiation processes, makes them explicit by phase-space integration and combines them with the virtual corrections. The resulting expressions read
@@
NNLOJET implements the antenna subtraction method [Gehrmann-DeRidder:2005btv, Daleo:2006xa, Currie:2013vh], which derives the subtraction terms starting from colour-ordered decompositions of all real-radiation contributions. The colour-ordering allows identifying each unresolved real-radiation configuration with a pair of hard radiator partons, which emit the unresolved partons. This configuration is then described by an antenna function consisting of the hard radiators and unresolved radiation in between them. Antenna functions were derived at NLO and NNLO for quark–antiquark [Gehrmann-DeRidder:2004ttg], quark–gluon [Gehrmann-DeRidder:2005svg] and gluon–gluon [Gehrmann-DeRidder:2005alt] pairs as hard radiators. All quarks are treated as massless. The unintegrated subtraction terms are obtained from products of antenna functions and reduced matrix elements of lower partonic multiplicity, which are summed over all configurations of hard radiators and unresolved partons. The kinematics of the antenna functions and the reduced matrix elements are factorized from each other through a phase-space mapping [Kosower:2002su, Daleo:2006xa], such that the antenna function can be integrated over the antenna phase space of the hard radiator partons and the unresolved radiation while leaving the kinematics of the reduced matrix element unaffected. The integrated antenna functions enter the subtraction terms and where they establish the explicit cancellation of infrared -poles between real and virtual contributions. They were computed up to NNLO for all different configurations of the hard radiators: final–final [Gehrmann-DeRidder:2005btv], initial–final [Daleo:2009yj] and initial–initial [Gehrmann:2011wi, Boughezal:2010mc, Gehrmann-DeRidder:2012too].
With these ingredients, all terms in equation (2) can be implemented in the antenna subtraction formalism, as described in detail in [Currie:2013vh]. For all processes available in NNLOJET, this implementation has been made on a process-by-process basis, usually starting from the implicit infrared singularity structure of the double-real-radiation process . A fully automated construction of the antenna subtraction terms in colour space is currently under development [Chen:2022ktf, Gehrmann:2023dxm].
@@ 3 The NNLOJET framework
The NNLOJET code is a parton-level event generator that provides the framework for the implementation of jet production processes to NNLO accuracy in QCD, using the antenna subtraction method. It contains the event generator infrastructure (Monte Carlo phase-space integration, event handling and analysis routines) and provides the unintegrated and integrated antenna functions and the phase-space mappings for all kinematic configurations. The NNLOJET phase-space integration is based on optimized parametrizations of the respective subprocess phase spaces (described in detail in [Gehrmann:2018szu]), which are then integrated using the adaptive Monte Carlo routine Vegas [Lepage:1977sw].
The implementation of processes in the NNLOJET framework requires the availability of the matrix elements for all RR, RV, and VV contributions, as well as the construction of the antenna subtraction terms. The subtraction terms have been validated by verifying their point-wise convergence to the respective subprocess matrix elements in all limits, as described in detail in [NigelGlover:2010kwr, Chen:2022ktf]. Some of the one-loop subprocess matrix elements were taken from the MCFM code [Campbell:2019dru], in particular for processes involving vector bosons [Campbell:2002tg]. The two-loop virtual corrections to all processes with four-point kinematics or beyond contain various special functions, in particular harmonic polylogarithms [Remiddi:1999ew] and their generalizations. These are evaluated using HPLOG [Gehrmann:2001pz], TDHPL [Gehrmann:2001jv] and CHAPLIN [Buehler:2011ev], which are all included as part of the NNLOJET distribution, with explicit consent of the respective code authors and thus in line with the underlying licensing provisions. The evaluation of parton distributions and of the strong coupling constant is performed using the standard LHAPDF [Buckley:2014ana] interface.
@@ 3.1 External dependencies
The NNLOJET code is mainly written in modern Fortran with some modules, dependencies and driver files written in C++ and Python. Multi-threading capabilities are supported through OpenMP. The NNLOJET installation requires compilers for Fortran/C/C++, a Python 3 interpreter (minimum version 3.10), as well as CMake (minimum version 3.18) to be available on the system. In addition, LHAPDF 6 [Buckley:2014ana] must be present, ideally with the lhapdf-config executable searchable from the $PATH of the system. The user is responsible for downloading the necessary PDF sets to be used in the calculation.
@@ 4 Processes in NNLOJET
The processes included in the current release are described in the following. Details on their NNLOJET implementation and on validation tests are described in detail in the respective publications. In general, the following checks have been applied on all process: (1) point-tests [NigelGlover:2010kwr, Chen:2022ktf] and (2) technical-cut checks for RR and RV contributions; (3) finiteness of the RV and VV contributions.
@@
The pole tests are performed numerically for a large number of bulk phase space points on the RV process, where -poles of matrix elements, mass-factorization terms and subtraction terms are implemented explicitly. In the VV process, all integrated subtraction terms are combined with the mass-factorization terms and the pole terms extracted [Catani:1998bh] from the double-virtual matrix element to obtain analytical pole cancellation and to define the finite remainder that is implemented numerically.
@@ 4.1 Jet production in electron–positron collisions
Jet final states in annihilation and related event shapes played a key role in establishing QCD as theory of the strong interaction and their theoretical study has been a major driver of developments in precision calculations in QCD. The cross section and related event shapes were the first jet observables to be computed to NNLO in QCD [Gehrmann-DeRidder:2007vsv], originally implemented in the code EERAD3 [Gehrmann-DeRidder:2014hxk] which is also based on antenna subtraction. The implementation of and in NNLOJET is described in [Gehrmann:2017xfb]. It uses the matrix elements for vector-boson decay into partons [Hagiwara:1988pp, Berends:1988yn, Bern:1997sc, Garland:2001tf, Garland:2002ak, Gehrmann:2002zr, Gehrmann:2011ab], which account for the incoming lepton kinematics. The commonly used LEP-era event-shape variables and jet algorithms [Biebel:2001dm] are part of the generic NNLOJET infrastructure.
@@ 4.2 Jet production in electron–proton collisions
Jet production in deeply inelastic electron–proton scattering (DIS) allows to probe aspects of the proton structure that are difficult to assess in inclusive DIS and provides important precision tests of QCD. The implementation of up to NNLO QCD is described in [Currie:2017tpe, Niehues:2018was], including both neutral-current and charged-current DIS. It is based on crossings [Gehrmann:2002zr] of the matrix elements to DIS kinematics. Jets can be defined in the Breit frame or in the electron–proton collider frame and DIS event shapes [Gehrmann:2019hwf] are part of the NNLOJET infrastructure.
@@ 4.3 Jet production in hadronic collisions
Jet production is the most ubiqutous QCD process at hadron colliders, probing the basic parton scattering reactions. A number of kinematic distributions can be derived from single-inclusive jet production and from dijet production . Their NNLOJET implementation is described in detail in [NigelGlover:2010kwr, Gehrmann-DeRidder:2011jwo, Gehrmann-DeRidder:2012dog, Currie:2016bfm, Currie:2017eqf, Gehrmann-DeRidder:2019ibf, Chen:2022tpk, Chen:2022clm]. It is based on the relevant subprocess matrix elements up to NNLO [Berends:1987me, Mangano:1990by, Kuijf:thesis, Bern:1993mq, Bern:1994fz, Kunszt:1994nq, Signer:thesis, Anastasiou:2001sv, Glover:2001af, Glover:2001rd, Anastasiou:2000kg, Anastasiou:2000ue, Anastasiou:2000mv]. The generic NNLOJET infrastructure allows selecting various different jet algorithms, performing scans over jet-resolution parameters and the setting of renormalization and factorization scales [Currie:2018xkj] either in a jet-based or event-based manner.
@@ 4.4 Vector-boson-plus-jet production
Precision studies of the electroweak interaction are commonly performed through massive vector-boson production, identified through their leptonic decay mode. Further kinematic information can be obtained by considering vector-boson-plus-jet final states. Observables derived from vector-boson and vector-boson-plus-jet production are widely used in precision measurements of electroweak parameters and of the proton structure. The NNLOJET implementation for , is described in [Gehrmann-DeRidder:2015wbt, Gehrmann-DeRidder:2016cdi, Gehrmann-DeRidder:2016jns, Gauld:2017tww, Gehrmann-DeRidder:2019avi, Gauld:2021pkr, Gehrmann-DeRidder:2023urf] and for , in [Gehrmann-DeRidder:2017mvr, Gehrmann-DeRidder:2019avi]. The implementations are based on parton-level matrix elements up to NNLO, obtained as kinematic crossings [Gehrmann:2011ab] of the matrix elements. Decays into lepton pairs and the off-shell vector-boson propagator are included. It is also possible to obtain simplified predictions for on-shell vector-boson production in the narrow-width prescription.